Determine whether each of the function from 𝑍 to Z is onto


(a) 𝑓(𝑛) = 𝑛³


(b) 𝑓(𝑛) = 𝑛² βˆ’ 1

1
Expert's answer
2022-03-21T00:32:45-0400

"a\,\,-\,\,no: f\left( n \right) =2\Rightarrow n^3=2-no\,\,solution\\b\,\,-\,\,no: f\left( n \right) =-2\Rightarrow n^2-1=-2-no\,\,solution"